Gareth owns 1 share of stock A and 1 share of stock B. In 1 year from today, the total value of his holdings is expected to be 122.28 dollars. Stock A is currently priced at 56.96 dollars, has an expected return of 17.82 percent, and is expected to pay a dividend of 3.2 dollars in 1 year from today. Stock B is currently priced at 54.76 dollars and is expected to pay a dividend of 5.65 dollars in 1 year from today. What is the expected return for stock B? Answer as a rate in decimal format so that 12.34% would be entered as .1234 and 0.98% would be entered as .0098.
Answer #1

Value of share = (Dividend in Year 1 + Price in Year 1)/(1+Expected Return)

Stock A

56.96 = (3.2+Price in year 1)/(1+17.82%)

Price in year 1 = 63.91

Value of Stock B after 1 year = 122.28 - 63.91 = $58.37

54.76 = (5.65+58.37)/(1+expected Return)

Expected return = 16.91%

i.e. 0.1691